there are two consecutive natural numbers,such that twice the square of the first is 34 more than the square of the second. find the numbers ??? i cannot get the answer of this question.
Answer by stanbon(75887)

there are two consecutive natural numbers,such that twice the square of the first is 34 more than the square of the second. find the numbers
------
1st: x
2nd: x+1
-----
Equation:
2x^2 - (x+1)^2 = 34
------
2x^2 - [x^2 + 2x + 1] = 34
------
2x^2 - x^2 -2x -1 = 34
-----
x^2 -2x -35 = 0
------
(x-7)(x+5) = 0
----
Answers:
x = 7
x+1 = 8
==============
Cheers,
Stan H.
==============